### v2

@@ 4 Simulations
[…]
In the main text, we report results for two different designs (Models 1 and 2) in which deviations from the null are such that only one of the $Y_{1},\ldots,Y_{p}$ violates independence. In Appendix C.1, we consider a third design (Model 3) in which all $Y_{1},\ldots,Y_{p}$ violate independence. The qualitative conclusions therein are similar to those in the main text.
Model 1:
Let $(X,Y_{1},\ldots,Y_{p})^{\prime}\sim\mathcal{N}(0,\Sigma_{\rho,\tau})$, where
$\Sigma_{\rho,\tau}=\begin{bmatrix}1&\rho&0&\ldots&0&\\ \rho&1&\tau&\ldots&\tau\\ 0&\tau&1&\ldots&\tau\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&\tau&\tau&\ldots&1\end{bmatrix}\in\mathbb{R}^{(1+p)\times(1+p)}.$
In this model, $H_{0}$ is satisfied when $\rho=0$. If $\rho\neq 0$, then $\rho$ determines the magnitude of the violation of the first hypothesis $H_{0,1}\colon Y_{1}\perp X$, while all other hypotheses $H_{0,j}\colon Y_{j}\perp X$, $j\geq 2$ hold. $\tau$ determines the correlation between $Y_{1},\ldots,Y_{p}$. [^2] This model is similar to Example 5.3 of Shi, Hallin, Drton, and Han (2022), except that we allow for all correlations between $Y_{1},\ldots,Y_{p}$ to be nonzero and consider one-dimensional $X$.
Model 2:
Let $X\sim Unif[-1,1]$ and $(\epsilon_{1},\ldots,\epsilon_{p})\sim\mathcal{N}(0,\Sigma_{\tau})$, where $\Sigma_{\tau}$ has diagonal elements equal to one and off-diagonal elements equal to $\tau$, and $(\epsilon_{1},\ldots,\epsilon_{p})$ is independent of $X$. Further, let
$Y_{j}=\left\{\begin{array}[]{cl}2\rho\cos(8\pi X)+\epsilon_{1},&\text{ for }j=1\\ \epsilon_{j}&\text{ for }2\leq j\leq p\end{array}\right..$
As in Model 1, $H_{0}$ is satisfied when $\rho=0$, and if $\rho\neq 0$, then $\rho$ determines the magnitude of the violation of the first hypothesis of independence. $\tau$ is the correlation between $Y_{1},\ldots,Y_{p}$. [^3] For $p=1$, this model is a reparameterized version of DGP (4) in Chatterjee (2021).
All the results in this section are for sample size $n=500$ and significance level $\alpha=0.05$.
4.1 Simulation I: Size and power for different block sizes
The first set of simulations concerns the rejection rates for different choices of the block size $q$. Figure 3 and 4 present the results for Models 1 and 2, respectively, based on $B=999$ bootstrap replications and $10,000$ Monte Carlo draws. First, we note that our baseline procedure BMB0 controls the size across different scenarios for all considered values of $q$, and its performance is very similar regardless of whether there is dependence between individual tests or not. It is, however, conservative in higher dimensions. The fact that the rejection rate is particularly low for $q=1$ is generally consistent with the upward bias in the bootstrap variance characterized in Lemma 1. The optimal rule from Section 2.3 yields $q^{*}(n)=3$ for $n=500$, marked by the vertical dashed lines in the graphs. This choice proves reasonable in all the considered scenarios.
@@ 4.1 Simulation I: Size and power for different block sizes
The issue of conservativeness is alleviated by the studentization. Both BMB1 and BMB2 effectively shrink the bootstrap distribution and yield rejection probabilities very close to the nominal level of 5% for small values of $q$. Since the correction in BMB1 is negligible for large $q$, the blue and black lines approach each other as $q$ increases. BMB2 maintains rejection rates closer to 5% as $q$ grows, but it generally slightly overrejects.
Panels C and D of Figure 3 indicate that all tests have very high power against the considered alternative in low dimensions under Model 1. The power generally deteriorates, albeit very slowly, as $p$ grows in this model. Apart from very small $q$ in BMB0, the power is not very sensitive to the choice of $q$. In Panels C and D of Figure 4, the simulated rejection rates are all equal to one, which is consistent with the observations that Chatterjee’s rank correlation is particularly well-suited for detecting oscillatory behavior (Chatterjee, 2021). As in the case of size, the power is not sensitive to correlation between $Y_{1},\ldots,Y_{p}$.
@@
Figure 3: Rejection rates for Model 1 under different choices of the big block size $q$.
Notes: The tests have nominal level of 5%. The null of independence holds if $\rho=0$. $\tau$ denotes the correlation between components of $Y$. Results for sample size $n=500$, $B=999$ bootstrap replications, and $S=10,000$ Monte Carlo draws.
Figure 4: Rejection rates for Model 2 under different choices of the big block size $q$.
@@
Notes: The tests have nominal level of 5%. The null of independence holds if $\rho=0$. $\tau$ denotes the correlation between components of $Y$ under the null. Results for sample size $n=500$, $B=999$ bootstrap replications, and $S=10,000$ Monte Carlo draws.
4.2 Simulation II: Power curves
@@
Since our test was seen in the previous subsection to be fairly insensitive to the particular choice of block size $q$, we now analyze power only for the optimal choice $q^{*}(n)=3$. We use $B=499$ bootstrap replications and $S=5,000$ Monte Carlo draws in this subsection.
We compare our test to the distance covariance test of Székely, Rizzo, and Bakirov (2007) and its adaption to high-dimensional settings proposed by Székely and Rizzo (2013). Specifically, we use the commands dcov.test and dcorT.test from the R package energy. In Appendix C.2, we also compare our test to tests proposed in Zhu, Zhang, Yao, and Shao (2020) and Zhou, Xu, Zhu, and Li (2024), but since these are computationally demanding, we are able to only run a small simulation experiment. In these, the additional tests perform very similarly to the distance covariance tests reported in the this section. Remember from the discussion in Section 1 that all of these alternative tests were developed as tests of the hypothesis $H_{0}^{joint}\colon(Y_{1},\ldots,Y_{p})\perp X$, which implies, but is not implied by, our hypothesis $H_{0}$. We are not aware of theoretical guarantees that these tests also control the rejection probability under $H_{0}$. The results in Figure 5 show, however, that at least in these simulations, the alternative tests approximately control the rejection frequency under $H_{0}$. Remarkably, the power curves for our test are very similar for both values of $\tau$ and the power decreases only slowly with $p$.
The power of any test of independence depends on the type of alternatives under consideration. Under Model 1 and when the dimension $p$ is not too large, the distance covariance tests exhibit higher power than ours. In high dimensions, however, our test dominates. One can also see that the distance covariance tests’ power decreases when the $Y_{1},\ldots,Y_{p}$ are dependent, while our tests’ power is not affected by this dependence. Under Model 2, the distance covariance tests have no power while our test remains powerful with power curves that are almost insensitive to the dimension $p$ and the degree of dependence among $Y_{1},\ldots,Y_{p}$.
Figure 5: Power curves of the BMB for Chatterjee’s rank correlation and distance covariance tests.
@@ 5 Empirical Application
We illustrate the practical usefulness of our proposed test by revisiting the study conducted by Hughes, DiTacchio, Hayes, Vollmers, Pulivarthy, Baggs, Panda, and Hogenesch (2009), which investigates transcriptional oscillations from mouse liver, NIH3T3, and U2OS cells. The goal is to identify genes whose transcript levels oscillate during the cell cycle. These cycles are crucial, as they play a key role in regulating metabolism and liver function, e.g., many liver genes operate on daily cycles, influencing vital processes such as detoxification, energy metabolism, and hormone regulation.
Hughes, DiTacchio, Hayes, Vollmers, Pulivarthy, Baggs, Panda, and Hogenesch (2009) collected liver tissue samples from mice at hourly intervals over a 48-hour period, pooling samples from 3-5 mice at each point. Their gene-level expression data set is available from Gene Expression Omnibus (GEO). In this section, we focus on the liver data set (accession GSE11923). We extracted the data using the GEOquery and BiocManager R Packages. The final data set contains $p=45,101$ genes. For each gene $j$, we observe $n=48$ transcript level measurements $Y_{j,1},\ldots,Y_{j,n}$ at different points in time, recorded in $X_{1},\ldots,X_{n}$.
@@
We apply our proposed bootstrap test based on the studentized test statistic BMB1 described in Remark 3, combined with the stepwise procedure developed in Section 3, both implemented via the R package hdIndep. We set $\alpha=0.05$, the nominal level at which the family-wise error rate is to be controlled, and the number of bootstrap samples to $B=1,000$. We choose the optimal block size developed in Section 2.3, which in this application is $q^{*}(n)=1$.
@@
The original study by Hughes, DiTacchio, Hayes, Vollmers, Pulivarthy, Baggs, Panda, and Hogenesch (2009) identified $3,667$ gene transcripts showing oscillatory behavior, but based on a different methodology for testing a different hypothesis than ours. [^4] The authors identify oscillatory patterns by testing for hidden periodicities in gene transcriptions using Fisher’s G and Straume (2004) tests. These tests are designed to test the null that the data are generated by a Gaussian white noise process against the alternative that the data is generated by a Gaussian white noise with a deterministic sinusoidal component. The authors then obtain so-called q-values following Storey and Tibshirani (2003) so as to select gene transcripts whose q-values are less than $\alpha=0.05$. Their procedure is guaranteed to control the false discovery rate at $\alpha=0.05$. Interestingly, our procedure finds more genes than the original study while at the same time providing stronger guarantees in the form of family-wise error rate control.
Among the genes identified by our method, $,91$ 919 were also identified in the original study, thus demonstrating substantial overlap. On the other hand, it singled out $2,635$ genes not identified in the original study (about $5.8\%$ of all gene transcripts), thus highlighting its ability to detect novel rhythmic gene activity. Figure 6 shows transcript level measurements of a random sample of six genes identified by our method, but not in the original study. A complete list of identified gene transcripts is available in the replication code (see replication vignette in accompanying R package hdIndep).
(a) Gene Symbol: Hdac5
(b) Gene Symbol: Lrrc3
(c) Gene Symbol: Mrps2
(d) Gene Symbol: Uap1
(e) Gene Symbol: Mir5136
(f) Gene Symbol: Impg1
Figure 6: Visualization transcript levels of 6 genes uniquely identified by our proposed stepdown method in Section 3, with $B=1000$ Gaussian multiplier bootstrap samples, $\alpha=0.05$, and optimal $q^{*}(n)=1$.

### v3

@@ 4 Simulations
In this section, we report results from a series of extensive simulation experiments in which we studied the finite-sample performance of our new test and compared it to existing tests.
4.1 Overview and Setup
First, we examine the influence of the choice of block size $q$ on the finite-sample performance of our test. Having established that the test’s rejection frequency is fairly insensitive to $q$, in subsequent experiments, we only consider our test with the optimal choice derived in Section 2.3. In the second set of experiments, we extensively study size control and power of our test. We consider a variety of data-generating processes. Key parameters that we vary in the simulations are the degree of dependence among $Y_{1},\ldots,Y_{p}$, the dimension $p$, and whether alternatives are sparse or dense. Having confirmed the theoretical findings that our test controls size and has power against all alternatives, we then show that existing tests based on distance covariance do not control size under our null hypothesis. They are valid only under the stronger null $H_{0}^{joint}$ mentioned in the introduction, i.e. when the copula of $Y_{1},\ldots,Y_{p}$ given $X$ does not depend on $X$. Finally, we compare our test to existing tests in the special case in which the copula of $Y_{1},\ldots,Y_{p}$ given $X$ does not depend on $X$, because these other tests are valid in that case.
4.1.1 Data-Generating Processes
Let $X$ be distributed uniformly on [1,1][-1,1], and let $(\epsilon_{1},\ldots,\epsilon_{p})\sim\mathcal{N}(0,\Sigma_{\tau})$ be a random vector independent of $X$, where $\Sigma_{\tau}\in\mathbb{R}^{p\times p}$ has diagonal elements equal to one and off-diagonal elements equal to $\tau$. We consider a range of deviations from the null of pairwise independence that differ in the functional form of the association between $X$ and elements of $Y_{1},\ldots,Y_{p}$, as well as in the number and strength of violations of individual hypotheses. In all considered models, the magnitude of the violations is parameterised by $\rho\in\mathbb{R}$, with the null hypothesis corresponding to $\rho=0$, while $\tau$ is the correlation between $Y_{1},\ldots,Y_{p}$ under the null.
Model 1 (Sparse linear alternatives).
$Y_{1}=\rho X+\varepsilon_{1}$ and $Y_{j}=\epsilon_{j}$ for $j\in\{2,\ldots,p\}$.
Model 2 (Sparse cosine alternatives).
$Y_{1}=\rho\cos(8\pi X)+\varepsilon_{1}$ and $Y_{j}=\epsilon_{j}$ for $j\in\{2,\ldots,p\}$.
Model 3 (Decaying linear alternatives).
$Y_{j}=0.9^{\,j-1}\rho X+\epsilon_{j}$ for $j\in\{1,\ldots,p\}$.
Model 4 (Decaying cosine alternatives).
$Y_{j}=0.9^{\,j-1}\rho\cos(8\pi X)+\epsilon_{j}$ for $j\in\{1,\ldots,p\}$.
Model 5 (Dense linear alternatives).
$Y_{j}=\rho X+\epsilon_{j}$ for $j\in\{1,\ldots,p\}$.
Model 6 (Dense cosine alternatives).
$Y_{j}=\rho\cos(8\pi X)+\epsilon_{j}$ for $j\in\{1,\ldots,p\}$.
In Models 1 and 2, $\rho$ determines the magnitude of the violation of the first hypothesis $H_{0,1}\colon Y_{1}\perp X$, while all other hypotheses, $H_{0,j}\colon Y_{j}\perp X$ for $j\in\{2,\ldots,p\}$, hold. In Models 5 and 6, all individual hypotheses are violated to the same degree, while Models 3 and 4 offer an intermediate scenario where the magnitude of the violations is strongest for the first hypothesis and decays exponentially with the index of the hypothesis. The individual violations are reparameterised versions of data-generating processes considered in the simulation study by Chatterjee (2021). Note that under the null, when $\rho=0$, all six models are identical.
We include cosine alternatives in these experiments partly because, in our empirical application, we conjecture such alternatives to be reasonable departures from independence.
4.1.2 Tests
We consider three variants of our proposed test that employ different types of studentisation. Implementations can be found in our accompanying R package hdIndep.
BMB0.
This test constructs the test statistic $\hat{T}$ based on the individual Chatterjee’s rank correlations $\hat{\xi}_{j}$ as in (2) and employs the critical value $\hat{c}(\alpha)$ from the distribution of the bootstrap statistic $\hat{T}^{B}$ as in (4).
BMB1.
This test is a variant of BMB0 that uses the studentised test statistic $\hat{T}^{stud}$ and the critical value from the bootstrap statistic $\hat{T}^{B,stud\,1}$ as in Remark 3.
BMB2.
This test is a variant of BMB0 that uses the studentised test statistic $\hat{T}^{stud}$ and the critical value from the bootstrap statistic $\hat{T}^{B,stud\,2}$ as in Remark 3.
For the special case in which the copula of $Y_{1},\ldots,Y_{p}$ given $X$ does not depend on $X$, we also compare our tests to the following alternatives:
dcov.test
The distance covariance test of Székely et al. (2007).
dcorT.test
The distance correlation t-test for high dimension proposed by Székely and Rizzo (2013).
ZXZL
Various rank-based tests proposed in Zhou et al. (2024): Hoeffding’s D (ZXZL-D), Blum-Kiefer-Rosenblatt’s R (ZXZL-R) and Bergsma-Dassios-Yanagimoto’s $\tau^{*}$ (ZXZL-$\tau^{*}$).
ZZYS-agg.dcov
The test based on aggregation of marginal distance covariances proposed by Zhu et al. (2020).
4.1.3 Implementation Details
All simulations are based on $B=499$ bootstrap replications, and $S=5000$ Monte Carlo draws, except in Experiment 4.2, where $S$ is reduced to 1000 due to mitigate computational burden. The significance level is set to $\alpha=0.05$ throughout. Additional results are reported in Appendix C and are omitted from the main text for brevity.
4.2 Experiment 1: BMB’s Insensitivity to Block Size
The first set of simulations concerns the rejection rates of our proposed test for different choices of the block size $q$. Figure 3 presents the results under the null in Panel A, and under the alternatives specified by Models 1–6 in Panels B–G. We focus on the case in which there is no correlation between $Y_{1},\ldots,Y_{p}$ under the null ($\tau=0$). The results are very similar when $\tau=0.5$; see Figure 10 in Appendix C.
@@ 4.2 Experiment 1: BMB’s Insensitivity to Block Size
First, we note that our baseline procedure BMB0 controls the size across different scenarios for all considered values of $q$. It is, however, conservative in higher dimensions. The fact that the rejection rate is particularly low for $q=1$ is consistent with the upward bias in the bootstrap variance characterised in Lemma 1. The optimal rule from Section 2.3 yields $q^{*}(n)=3$ for $n=500$, marked by the vertical dashed lines in the graphs. This choice proves reasonable in all the considered scenarios. The issue of conservativeness is alleviated by the studentisation. Both BMB1 and BMB2 effectively shrink the bootstrap distribution and yield rejection probabilities very close to the nominal level of 5% for small values of $q$. Since the correction in BMB1 is negligible for large $q$, the blue and black lines approach each other as $q$ increases. BMB2 maintains rejection rates closer to 5% as $q$ grows, but it generally slightly overrejects.
Panels B–G of Figure 3 indicate that all tests have high power in low dimensions. Under sparse and decaying alternatives, the power decreases as $p$ grows, while the power increases for dense alternatives.
@@
Figure 3: Rejection rates in Models 1–6 with $\tau=0$ under different choices of the big block size $q$.
@@
Notes: The tests have nominal level of 5%. The null of independence holds if $\rho=0$, while $\tau$ denotes the correlation between variables $Y_{1},\ldots,Y_{p}$ under the null. Results for sample size $n=500$, $B=499$ bootstrap replications, and $S=5,000$ Monte Carlo draws.
4.3 Experiment 2: BMB’s Size Control and Power
@@
Since, in Experiment 1, our test was seen to be fairly insensitive to the particular choice of block size $q$, we now analyse size control and power only for the optimal choice, which for $n=500$ is $q^{*}(n)=3$.
Figures 4 and 5 present the power curves of our test under linear and cosine alternatives, respectively. For each model, we consider the cases in which $Y_{1},\ldots,Y_{p}$ are mutually independent ($\tau=0$) or dependent ($\tau=0.5$). In each graph, we show rejection frequencies of our test as we vary $\rho$. The null hypothesis corresponds to $\rho=0$, and as $\rho$ increases the violation of the null becomes larger.
Figure 4: Power curves under linear alternatives in high dimensions.
@@
Figure 5: Power curves under cosine alternatives in high dimensions.
First, the three variants of our test, BMB0, BMB1, and BMB2, control size across scenarios, including high-dimensional settings (large $p$) and those with dependence among $Y_{1},\ldots,Y_{p}$ ($\tau=0.5$).
Second, our test has power against all considered alternatives. Interestingly, comparing the power curves for $\tau=0$ and $\tau=0.5$, we see that the power of our test is not or only slightly affected by dependence among $Y_{1},\ldots,Y_{p}$.
In Appendix C, we present analogous results for sample sizes $n=200$ and $n=1000$. While the power of the test increases with sample sizes, the qualitative conclusions remain the same.
4.4 Experiment 3: Invalidity of Distance Covariance Tests
As indicated above, the validity of existing tests dcorT.test, dcov.test, ZXZL, and ZZYS-agg.dcov have been established under the stronger null $H_{0}^{joint}$ that $(Y_{1},\ldots,Y_{p})\perp X$. We now investigate the size control properties of these tests when our null hypothesis $H_{0}\colon Y_{j}\perp X$ for $j=1,\ldots,p$ holds, but $H_{0}^{joint}$ is violated.
To this end we consider a modified version of the data-generating processes Models 1–6, which are identical under the null, in which the dependence parameter $\tau$ depends on $X$. Specifically, we consider Model 1 with $\rho=0$ and instead of $(\epsilon_{1},\ldots,\epsilon_{p})\sim\mathcal{N}(0,\Sigma_{\tau})$, we now consider $(\epsilon_{1},\ldots,\epsilon_{p})|X\sim\mathcal{N}(0,\Sigma_{\tau}(X))$, where $\Sigma_{\tau}(X)$ has diagonal elements equal to one and off-diagonal elements equal to $\tau(X):=\gamma(1+X)/2$. Since $\rho=0$, the marginal distributions of $Y_{j}$ are all independent of $X$, but the correlation parameter among $Y_{1},\ldots,Y_{p}$ is a function of $X$. The parameter $\gamma$ governs the strength of that dependence.
Figure 6 presents the rejection rates of all tests under the null hypothesis $H_{0}$ when $p=50$. Our test controls size across all values of $\gamma$, while the distance covariance tests dcorT.test and dcov.test do not and their rejection rates increase with $\gamma$. Interestingly, the ZXZL and ZZYS-agg.dcov tests also control size. Since they are based on marginal distance covariance and rank-based statistics, we conjecture that these tests are not only valid under $H_{0}^{joint}$, but also under $H_{0}$.
Figure 6: Rejection rates under the null when the copula of $Y_{1},\ldots,Y_{p}$ given $X$ is not independent of $X$.
4.5 Experiment 4: Comparison to Other Tests
In this section, we compare our test to all other tests in the special case in which the copula of $Y_{1},\ldots,Y_{p}$ given $X$ does not depend on $X$, both under the null and the alternative hypotheses. We divide this experiment into two sub-experiments because the ZXZL and ZZYS-agg.dcov tests are computationally so demanding that we were not able to run experiments in high dimensional settings ($p$ large). Therefore, we first consider an experiment in which we compare our test to the distance covariance tests dcorT.test and dcov.test for a range of values of $p$ up to very high dimensions. In a second experiment, we then compare our test to the ZXZL and ZZYS-agg.dcov tests in more moderate dimensions up to $p=500$.
4.5.1 Experiment 4.1: Comparisons Including Very High-Dimensional Settings
Figures 4 and 5 present the power curves of our test under linear and cosine alternatives, respectively. For each model, we consider the cases in which $Y_{1},\ldots,Y_{p}$ are mutually independent ($\tau=0$) or dependent ($\tau=0.5$). In each graph, we show rejection frequencies of our test as we vary $\rho$. The null hypothesis corresponds to $\rho=0$, and as $\rho$ increases the violation of the null becomes larger.
Several interesting findings emerge from these figures. First, while our test is almost insensitive to dependence among $Y_{1},\ldots,Y_{p}$, the distance covariance tests dcorT.test and dcov.test are not. For instance, comparing Panel A and B in Figure 4 for sparse alternatives, we see that their power may suffer substantially when $Y_{1},\ldots,Y_{p}$ are dependent. A similar but less extreme pattern is observed under decaying linear alternatives (Panels C and D in Figure 4). Under dense alternatives, under which these tests are known to perform particularly well, their power is less affected by dependence among $Y_{1},\ldots,Y_{p}$.
Second, as expected, the distance covariance tests perform particularly well under dense alternatives, dominating the power curves of our tests. Our test performs particularly well under sparse and decaying alternatives, where it tends to be more powerful than the distance covariance tests, particularly so in high dimensions.
Third, considering the cosine alternatives (Figure 5), we see that our test is more powerful than the distance covariance tests in all scenarios, including even the dense alternatives in which one would expect the distance covariance tests to perform well.
In Appendix C, we present analogous results for sample sizes $n=200$ and $n=1000$. While the power of all tests increases with sample sizes, the qualitative conclusions remain the same.
4.5.2 Experiment 4.2: Comparisons Limited to Moderately High Dimensions
In this section, we compare our test to the ZXZL and ZZYS-agg.dcov tests. [^2] The tests ZXZL-D, ZXZL-R, ZXZL-$\tau^{*}$, and ZZYS-agg.dcov were run using the implementation of Zhou et al. (2024) provided at https://github.com/Yeqing-TJ/Rank-based-test-in-high-dimension [Accessed on March 10, 2025]. Due to the high computational complexity of these additional tests, we limit the simulation setup to samples of size $n=200$ with the maximum of $p=200$ individual hypotheses. [^3] Estimation of the variance of the rank-based indices of Zhou et al. (2024) in one sample with $p=500$ and $n=500$ takes over 40 minutes using an Intel i7-1185G7 @ 3.00 GHz processor, which renders simulations for such settings infeasible. For comparison, our bootstrap test with $B=499$ takes less than a second to compute in this setting.
Figures 7 and 8 show power comparisons analogous to Figures 4 and 5, respectively. The broad patterns are similar to those in Figures 4 and 5 in that the ZXZL and ZZYS-agg.dcov tests exhibit similar performance to the distance covariance tests studied in Experiment 4.1.
In the low-dimensional settings with linear alternatives (Figure 7), the ZXZL and ZZYS-agg.dcov tests dominate our test when $p$ is small, there is no dependence among $Y_{1},\ldots,Y_{p}$, or the alternatives are dense. Our test performs relatively better the larger $p$ and dominates the other tests when $p$ is sufficiently large (except for the dense alternatives). As in Experiment 4.1, our test’s power curves are barely affected by dependence among $Y_{1},\ldots,Y_{p}$, but the alternative tests’ power curves decrease substantially when $Y_{1},\ldots,Y_{p}$ are dependent; compare for instance Panels A and B in Figure 7.
Under cosine alternatives (Figure 8), the ZXZL and ZZYS-agg.dcov tests have virtually no power, and our test is more powerful than the competitors in all scenarios.
Figure 7: Power curves under linear alternatives in moderately high dimensions.
Notes: The tests have nominal level of 5%. Results for $n=200$, $B=499$ bootstrap samples, and $S=1,000$ Monte Carlo draws.
Figure 8: Power curves under cosine alternatives in moderately high dimensions.
Notes: The tests have nominal level of 5%. Results for $n=200$, $B=499$ bootstrap samples, and $S=1,000$ Monte Carlo draws.
4.6 Summary of Findings
The simulation experiments reveal several interesting findings:
  1. 1.
    Our test is fairly insensitive to the choice of block size $q$.
  2. 2.
    Our test’s performance is barely affected by the strength of dependence among $Y_{1},\ldots,Y_{p}$, while the alternative tests’ power deteriorates under dependence.
  3. 3.
    Our test is relatively more powerful in higher dimensions and dominates the other tests when $p$ is sufficiently large, except when alternatives are dense.
  4. 4.
    Our test is powerful against cosine alternatives, while the alternative tests are not.
  5. 5.
    The distance covariance tests dcorT.test and dcov.test may not control size under the null hypothesis $H_{0}$.
  6. 6.
    We did not find any evidence against validity of the ZXZL and ZZYS-agg.dcov tests under the null hypothesis $H_{0}$, but these tests are computationally much more demanding than ours and infeasible to simulate in high dimensions.
@@ 5 Empirical Application
We illustrate the practical usefulness of our proposed test by revisiting the study conducted by Hughes et al. (2009), which investigates transcriptional oscillations from mouse liver, NIH3T3, and U2OS cells. The goal is to identify genes whose transcript levels oscillate during the cell cycle. These cycles are crucial as they play a key role in regulating metabolism and liver function, e.g., many liver genes operate on daily cycles, influencing vital processes such as detoxification, energy metabolism, and hormone regulation.
Hughes et al. (2009) collected liver tissue samples from mice at hourly intervals over a 48-hour period, pooling samples from 3–5 mice at each point. Their gene-level expression data set is available from Gene Expression Omnibus (GEO). In this section, we focus on the liver data set (accession GSE11923). We extracted the data using the GEOquery and BiocManager R packages. The final data set contains $p=45,101$ genes. For each gene $j$, we observe $n=48$ transcript level measurements $Y_{j,1},\ldots,Y_{j,n}$ at different points in time, recorded in $X_{1},\ldots,X_{n}$.
@@
We apply our proposed bootstrap test based on the studentised test statistic BMB1 described in Remark 3, combined with the stepwise procedure developed in Section 3, both implemented via the R package hdIndep. We set $\alpha=0.05$, the nominal level at which the family-wise error rate is to be controlled, and the number of bootstrap samples to $B=1,000$. We choose the optimal block size developed in Section 2.3, which in this application is $q^{*}(n)=1$.
@@
The original study by Hughes et al. (2009) identified $3,667$ gene transcripts showing oscillatory behaviour, but based on a different methodology for testing a different hypothesis than ours. [^4] The authors identify oscillatory patterns by testing for hidden periodicities in gene transcriptions using Fisher’s G and Straume (2004) tests. These tests are designed to test the null that the data are generated by a Gaussian white noise process against the alternative that the data is generated by a Gaussian white noise with a deterministic sinusoidal component. The authors then obtain so-called q-values following Storey and Tibshirani (2003) so as to select gene transcripts whose q-values are less than $\alpha=0.05$. Their procedure is guaranteed to control the false discovery rate at $\alpha=0.05$. Interestingly, our procedure finds more genes than the original study while at the same time providing stronger guarantees in the form of family-wise error rate control.
Among the genes identified by our method, $1,919$ were also identified in the original study, thus demonstrating substantial overlap. On the other hand, it singled out $2,635$ genes not identified in the original study (about $5.8\%$ of all gene transcripts), thus highlighting its ability to detect novel rhythmic gene activity. Figure 9 shows transcript level measurements of a random sample of six genes identified by our method, but not in the original study. A complete list of identified gene transcripts is available in the replication code (see replication vignette in accompanying R package hdIndep).
Figure 9: Visualisation of transcript levels of six genes uniquely identified by our proposed stepdown method in Section 3, with $B=1000$ Gaussian multiplier bootstrap samples, $\alpha=0.05$, and optimal $q^{*}(n)=1$. Gene symbols: (a) Hdac5, (b) Lrrc3, (c) Mrps2, (d) Uap1, (e) Mir5136, (f) Impg1.